OpenStudy (anonymous):

can someone show me the steps to solving this equation 5/13 t = -9

OpenStudy (anonymous):

t = -117/5 or (5 t)/13+9 = 0

jhonyy9 (jhonyy9):

5t= - 117 t= -117/5